OpenStudy (anonymous):

Functions f(x) and g(x) are described as follows: f(x) = −5x2 + 9 x g(x) 0 0 1 4 2 8 3 4 4 0 Which statement best compares the maximum value of the two functions? It is equal for both functions. It is 3 units higher for f(x) than g(x). It is 3 units lower for f(x) than g(x). It is 1 unit higher for f(x) than g(x).

OpenStudy (freckles):

Well have you found the max values of both functions?

OpenStudy (freckles):

f is already in vertex form

OpenStudy (anonymous):

no i havent.

OpenStudy (freckles):

\[f(x)=a(x-h)^2+k \\ \text{ if } a \text{ is negative then } k \text{ is a \max} \\ \text{ if } a \text{ is positive then } k \text{ is a \min}\]

OpenStudy (freckles):

And for g you just have to look for which g(x) value is the highest and that is the max

OpenStudy (anonymous):

I am very confused i am just learning this stuff.

OpenStudy (freckles):

compare the following two things: \[f(x)=a(x-h)^2+k \\ f(x)-5(x-0)^2+9\] is there anyway you can determine what k is

OpenStudy (freckles):

everything is lined up where it should be a is -5 h is 0 k is...

OpenStudy (anonymous):

9?

OpenStudy (freckles):

yes

OpenStudy (freckles):

now followed what I said above

OpenStudy (freckles):

since a is negative then k is a max

OpenStudy (freckles):

you did have a is -5 so k=9 is your max

OpenStudy (freckles):

9 is the max value of f

OpenStudy (freckles):

can you look at your list of g(x) values and just say what the highest number is listed there

OpenStudy (anonymous):

8?

OpenStudy (freckles):

yes so the max of g is 8

OpenStudy (freckles):

are you able to answer the question now?

OpenStudy (anonymous):

SO the answer is c?

OpenStudy (freckles):

so you think 9 is 3 units less than 8?

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

Its D.

OpenStudy (freckles):

yes 9 is one more than 8

OpenStudy (freckles):

if you don't know the vertex form of a parabola you could graph the function given to you next time and just go to the peak of the mountain and find the highest y value there

OpenStudy (anonymous):

Can you help with 1 more?

OpenStudy (freckles):

I can try

OpenStudy (anonymous):

Mary wants to hang a mirror in her room. The mirror and frame must have an area of 7 square feet. The mirror is 2 feet wide and 3 feet long. Which quadratic equation can be used to determine the thickness of the frame, x? Square with an inner frame with height of 2ft on the left frame and width of 3ft on the top. Arrow on the bottom frame with an x and an arrow on the right frame with an x. x2 + 14x − 2 = 0 2x2 + 10x − 7 = 0 3x2 + 12x − 7 = 0 4x2 + 10x − 1 = 0

OpenStudy (anonymous):

let me get the graph

OpenStudy (anonymous):

http://prntscr.com/87afn6

OpenStudy (freckles):